Expert Q&A Done Determine the othewing 202.000.00 100 10023 D1
Answer #1

2.02 : 3 sig fig

8.302 : 4 sig fig

0.0297 : 3 sig fig

Now,

2.02 have 2 sig fig to the right of deciaml and 8.302 have 3 sig fig to the right of decimal. Thus, their sum must also have 2 sig fig to the right of the decimal.

2.02 + 8.302 = 10.322 = 10.32

Now,

10.32 have 2 sig fig to the right of decimal and 0.0297 have 3 sig fig to the right of decimal. Thus, their sum must also have 2 sig fig to the right of the decimal.

10.32 + 0.0297 = 10.3497 = 10.35

Thus,

2.02 + 8.302 + 0.0297 = 10.35